Circular No. 8804

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2007N–2007U Eight apparent supernovae have been discovered on unfiltered CCD

frames and reported to the Central Bureau: 2007N, 2007O, 2007P, and 2007T by W. Li (LOSS/KAIT; cf. IAUC 8798; 2007N and 2007O with N. Lee, 2007P with D. Winslow, and 2007T with D. R. Madison), and the rest by T. Puckett (0.35-m reflector at Ellijay, GA, except 0.40-m reflector at Portal, AZ, for 2007U; cf. IAUC 8798; 2007Q and 2007U with J. Newton, 2007R with P. Gray, and 2007S with R. Gorelli).

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007N Jan. 21.52 12^{h}49^{m}01.^{s}25 − 9^{o}27′10′′.2 17.3 4′′.3 W, 14′′.3 N
2007O Jan. 21.60 14 56 05.18 +45 24 16.9 16.8 8′′.2 E, 1′′.5 S
2007P Jan. 23.40 9 58 56.41 −18 49 11.4 18.0 4′′.8 W, 4′′.8 S
2007Q Jan. 3.42 15 13 08.07 +41 15 40.1 17.8 14′′.5 E, 13′′.8 S
2007R Jan. 26.10 7 46 37.53 +44 47 22.1 16.9 1′′.9 W, 3′′.9 S
2007S Jan. 29.24 10 00 31.26 + 4 24 26.2 16.7 11′′.4 W, 0′′.1 N
2007T Feb. 3.57 15 00 45.73 +49 59 24.3 16.9 3′′.5 W, 13′′.2 N
2007U Feb. 5.09 5 05 07.13 −18 59 19.9 17.5 16′′.4 W, 8′′.8 S

KAIT limiting magnitude for 2007N in MCG –01-33-12: 2006 Dec. 30.54 UT, [19.0. KAIT limiting magnitude for 2007O in UGC 9612: 2006 July 5.23, [19.5. Both 2007N and 2007O are type-Ia supernovae, discovered near maximum (spectroscopic details on CBET 818). Additional KAIT magnitudes for 2007P in ESO 566-G36: 2006 Feb. 2.38, [19.0; 2007 Jan. 13.41, [18.1; 24.38, 17.9. SN 2007P is probably a young type-II supernova (spectroscopic details on CBET 822). Additional magnitudes by Puckett for 2007Q in NGC 5888: 2005 Nov. 15, [19.0; 2007 Jan. 26.40, 18.1. SN 2007Q is a type-II supernova discovered near the time of explosion (spectroscopic details on CBET 840). Additional magnitudes for 2007R in UGC 4008: 2005 Dec. 25, [19.0 (Puckett); 2007 Jan. 27.12, 16.5 (T. Orff). SN 2007R is a type-Ia supernova next maximum (details on CBET 826). Additional magnitudes for 2007S in UGC 5378: 2005 Dec. 29, [19.0 (Puckett); 2007 Jan. 31.20, 16.6 (Orff). SN 2007S is a young type-Ia supernova (spectroscopic details on CBETs 829 and 839). Additional approximate KAIT magnitudes for 2007T in NGC 5828: 2006 Aug. 14.20, [19.0; 2007 Feb. 4.50, 16.8. SN 2007T is a type-II supernova, discovered near the time of explosion (details on CBET 837). Additional magnitudes for 2007U in ESO 552-65: Jan. 11, [19.0 (Puckett); Feb. 6.02, 17.5 (Orff). SN 2007U is a type-II supernova found some 10 days past explosion (spectroscopic details on CBET 838).

2007 February 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green